UT 20161022 LUCI1 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt
Observer AO: GTayler
Telescope Operator: GBechetti

Plan for the night

ARGOIS-team starts, we take over at 2.30am

- Stability and robustness of FLAO-LUCI wo AFC in Bin = 1 and 2 again

Summary

- Night with mostly brilliant seeing

- no progress to many disturbances...

- three different AO issues (lense, arbitrator, unknown flux)

- fire alarm at LBT and frozen obs1

Details

All times are UT
Files are luci1.20161022.NNNN.fits

00:50 Sunset, Weather: photometric, T=9deg, Hum=30%, low wind

about 8:00 Telescope got stuck

09:30 Everything seems back alive now, Genos powers up what he need, Greg restarts the AdSec

09:40 Geno checks pointing and collimates, seeing on guider 0.5", latest DIMM about 2hrs ago 0.6".

Robustness & stability of FLAO-LUCI1

Bin=2

09:48 Preset to T Tauri, TTau_H2_rev.xml, AOref has 9.2 in Nomad, 10.6 in FLAO, RunAO failed, camera lense did not converge,
resent preset, RunAO, OK, use ABS X = -7.5 next time for PA=-90,

images #0562-0572

Preset again, modified scipt with ABS X = -7.5, RunAO OK, first offset failed, issue with ADsec arbitrator,
preset again, RunAO, OK, images #0576-0599 are individual 2.6sec images and must be integrated,
star is brighter, ie Bin=1 was used not clear if this and to which amount affects PSF-check for TTAu (which is bin=2)

images #0573-0609, note: this are date fro HD27560 which is a nearby PSF-calibrator for TTAu

10:51 Preset back to TTAu, RunAO, OK, T Tau Sa can easily, T Tau Sb just be seen (wo any image processing) ,
loop opened at #0624, skip frame, reason unknown (probably due to frozen obs1?)

11:16 Fire alarm at LBT

11:19 We have to wait now, LBT people are out for fire alarm and Geno must boot obs1 (got frozen before) etc. before we can resume

11:26 They are back, Geno reboots obs1 + brings up lots of other things, reason for fire alarm: humidifier,

images #0613-0624

Bin=1

12:10 Preset to Trapezium, RunAO OK, LUCI GUI frozen, large offsets OK, 2nd offset in patrol field failed, star wrong flux,
resent preset, LUCI GUI back alive, RunAO OK, LUCI GUI freezes again, keep observing until
loop opens due to bright sky (sun rises in 30mins)

images #0625-660

13:05 Loop opened as background is too high, stopped, end of tests

dimm.20161022jhill.jpg

-- %USERSIG{JochenHeidt - 2016-10-22}%
Topic revision: r4 - 24 Oct 2016, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback